I need to find the slope of a line parallel to each line.
1.y=2x-3 2.8y=-7x+16

Answer by checkley71(8403)

The slope for the first equation is 2. using the standard form y=mx+b where the m=slope.
The qecond equation needs to be changed to the standard slope form thus
2.8y=-7x+18 or y=(-7x+18)/2.8 or y=-7/2.8x+18/2.8 or y=-2.5+6.43 or this slope is -2.5.
Any line with a slope of 2 or -2.5 is parallel to these lines.